OpenStudy (anonymous):

Hi guys! I don't understand why the total angular momentum of deuterium is l(/aɪ/)=1. What are the values for total spin S and L? thanks

OpenStudy (experimentx):

the angular momentum of deuterium is = angular momentum of 1 protons + 1 neutron

OpenStudy (experimentx):

for proton: spin = 1/2 for neutron: spin = 1/2 I guess the orbital quantum number is zero for both

OpenStudy (anonymous):

ok but if I told you that the total angular momentum I=1 and S=1 so for L I can write: \[|I-S|\le L \ge L+S\] and so L can asuume the values 0,1,2 but the teory says that L=1 is forbidden by violation of parity! the question is: how can I get L=2?